# Efficient Reasoning

### v1

@@ Abstract
Fine-tuning large language models (LLMs) on chain-of-thought (CoT) data shows that a small amount of high-quality data can outperform massive datasets. Yet, what constitutes “quality” remains ill-defined. Existing reasoning methods rely on indirect heuristics such as problem difficulty or trace length, while instruction-tuning has explored a broader range of automated selection strategies—but rarely in the context of reasoning. We propose to define reasoning data quality using influence functions, which measure the causal effect of individual CoT examples on downstream accuracy, and introduce influence-based pruning, which consistently outperforms perplexity and embedding-based baselines on math reasoning within a model family.
@@ 1 Introduction
Large language models (LLMs) show strong reasoning gains when fine-tuned on chain-of-thought (CoT) data (DeepSeek-AI and Others,, 2025). Recent studies suggest that data quality often outweighs quantity (Rodchenko et al.,, 2025; Yuan et al.,, 2025; Ye et al.,, 2025), yet what constitutes “quality” remains unclear.
Prior work on reasoning has used proxies for data quality, such as question difficulty (Chen et al.,, 2025) or CoT length (Yuan et al.,, 2025), which offer only indirect signals of downstream performance. In contrast, instruction-tuning studies have systematically evaluated automated data selection methods—perplexity filtering, embedding similarity, gradient-based scoring (Yin and Rush,, 2025; Ivison et al.,, 2025)—but these have not been fully applied to reasoning data.
@@ 2 Related Works
Heuristic methods define scores $s(d)$ directly from inherent properties of each training example $d\in\mathcal{D}$, independent of $\mathcal{V}$ or a specific model. Recent works take into account heuristics such as estimated question difficulty, CoT length, and problem diversity (Ye et al.,, 2025; Muennighoff et al.,, 2025). LIMO (Ye et al.,, 2025) also incorporates human evaluation, selecting solutions that exhibit desirable behaviors such as elaboration and self-verification. Select2Reason (Yang et al.,, 2025) provides a systematic study, showing that higher $s(d)$ assigned to examples with longer traces and harder problems yields stronger downstream performance, whereas diversity has limited benefit.
@@
These methods score each example $s(d)$ by its negative log-likelihood under the base model. Top-PPL selects high-perplexity (hard) cases, while Mid-PPL favors mid-range examples to avoid trivial or outlier data. These scores are cheap to compute and have shown to improve sample efficiency in the instruction tuning setting (Yin and Rush,, 2025; Ankner et al.,, 2024).
@@
Embedding-based methods use data representations to assign scores. S1 (Muennighoff et al.,, 2025) enforces coverage by classifying problems into topics with an LLM and sampling across them. RDS+ (Ivison et al.,, 2025) computes pairwise scores $s(v,d)=\langle E(v),E(d)\rangle$ between validation and training examples, selecting examples most relavent to each $v\in\mathcal{V}$ in a round-robin fashion.
@@
Gradient and influence based methods compute $s(v,d)$ by estimating the causal contribution of training examples on validation loss. TracIn (Pruthi et al.,, 2020) sums gradient inner products across training checkpoints and LESS (Xia et al.,, 2024) scales this approach to large models using gradient datastores. InfDist (Nikdan et al.,, 2025) learns optimal per-example weights so that a single gradient step minimizes validation loss.
@@
Unlike gradient-similarity methods that track training dynamics, influence functions measure the effect of a training example on converged model parameters and downstream objectives $f(\theta^{\star})$. Formally, the influence of a training example $d\in\mathcal{D}$ is $\mathcal{I}_{f}(d)=-\nabla_{\theta}f(\theta^{\star})^{\top}H^{-1}\nabla_{\theta}L(d,\theta^{\star})$ where $H$ is the Hessian of the training objective at $\theta^{\star}$. Approximations such as EK-FAC have made this tractable for LLMs (Grosse et al.,, 2023). Ruis et al., (2025) shows that examples with high estimated influence on cross-entropy loss also causally affect downstream accuracy, and their removal causes larger performance drops than random pruning or TracIn.
@@ 3.1 Influence Functions Scoring
Let $\theta$ be a pretrained model and $\mathcal{D}$ the LIMO training dataset (Ye et al.,, 2025). We fine-tune $\theta$ on $\mathcal{D}$ for 10 epochs following the settings of Ye et al., (2025), obtaining $\theta^{\text{LIMO}}$. We then evaluate on the MATH500 benchmark $\mathcal{V}$ for both model parameters $\theta$ and $\theta^{\text{LIMO}}$ (Hendrycks et al.,, 2021).
@@
Our goal is to estimate the causal influence of training examples $d\in\mathcal{D}$ on these outcome sets. Because influence functions require a differentiable proxy, we follow Grosse et al., (2023) and compute influence with respect to cross-entropy loss on completions. For each validation query $v\in C\cup I$ with completion $y_{c}$, the influence of a training point $d$ is:
@@ 4 Experiments and Results
We perform supervised fine-tuning (SFT) on pruned versions of the LIMO dataset, using two models: LLaMA-3-8B-Instruct (Grattafiori et al.,, 2024), also used during pruning, and Qwen2.5-Math-7B-Instruct (Yang et al.,, 2024), external to the pruning process. We ran full finetuning following the training settings from LIMO for 10 epochs, saving checkpoints at every other epoch. Given that we already were optimizing for MATH500, we treated it as a validation set and chose the checkpoint which performed best on MATH500. We chose this strategy as for different datasets, optimal checkpoints were not consistent.
Evaluation focuses on the math reasoning benchmarks: AIME24, AMC23, OlympiadBench (He et al.,, 2024), GSM8k (Cobbe et al.,, 2021). Additionally we include GPQA (Rein et al.,, 2024) to assess transfer to non-math reasoning. We compare three IF-based pruning strategies (Correct, Incorrect, Combined) against Random, Mid-PPL , and RDS+ (Ivison et al.,, 2025), and report also the unpruned dataset and the base model for reference. All results use the unbiased pass@1 metric (Chen et al.,, 2021) with N=8N{=}8 samples in a zero-shot chain-of-thought setting, following the LIMO protocol. Full training and evaluation details are provided in the Appendix.
@@ Appendix A Training and Evaluation Details
We follow the evaluation setup of LIMO (Ye et al.,, 2025), with the addition of the GSM8k test set manually added. We report the unbiased pass@1 metric (Chen et al.,, 2021), estimated from N=8N{=}8 sampled generations. Sampling uses temperature $0.6$, top-$p=0.95$, top-$k=1$, and a maximum generation length of 16,38416{,}384 tokens.
@@ Appendix B Notes on influence functions computation
To compute influence function scores, we follow the derivation and approximations described in Grosse et al., (2023). Following Ruis et al., (2025),we perform the computation considering only MLP layers.

### v2

@@ Abstract
Fine-tuning large language models (LLMs) on chain-of-thought (CoT) data shows that a small amount of high-quality data can outperform massive datasets. Yet, what constitutes “quality” remains ill-defined. Existing reasoning methods rely on indirect heuristics such as problem difficulty or trace length, while instruction-tuning has explored a broader range of automated selection strategies—but rarely in the context of reasoning. We propose to define reasoning data quality using influence functions, which measure the causal effect of individual CoT examples on downstream accuracy, and introduce influence-based pruning, which consistently outperforms perplexity and embedding-based baselines on math reasoning within a model family.  [^1] Code available at: https://github.com/mila-capitalone-collab/inf-functions-reasoning/tree/main
@@ 1 Introduction
Large language models (LLMs) show strong reasoning gains when fine-tuned on chain-of-thought (CoT) data (deepseekai2025deepseekr1incentivizingreasoningcapability). Recent studies suggest that data quality often outweighs quantity (rodchenko2025aiproblemdataproblem; yuan2025naturalreasoningreasoningwild28m; ye2025limoreasoning), yet what constitutes “quality” remains unclear.
Prior work on reasoning has used proxies for data quality, such as question difficulty (chen2025advancingmathematicalreasoninglanguage) or CoT length (yuan2025naturalreasoningreasoningwild28m), which offer only indirect signals of downstream performance. In contrast, instruction-tuning studies have systematically evaluated automated data selection methods—perplexity filtering, embedding similarity, gradient-based scoring (yin2025computeconstraineddataselection; rds)—but these have not been fully applied to reasoning data.
@@ 2 Related Works
Heuristic methods define scores $s(d)$ directly from inherent properties of each training example $d\in\mathcal{D}$, independent of $\mathcal{V}$ or a specific model. Recent works take into account heuristics such as estimated question difficulty, CoT length, and problem diversity (ye2025limoreasoning; s1). LIMO (ye2025limoreasoning) also incorporates human evaluation, selecting solutions that exhibit desirable behaviors such as elaboration and self-verification. Select2Reason (select2reason) provides a systematic study, showing that higher $s(d)$ assigned to examples with longer traces and harder problems yields stronger downstream performance, whereas diversity has limited benefit.
@@
These methods score each example $s(d)$ by its negative log-likelihood under the base model. Top-PPL selects high-perplexity (hard) cases, while Mid-PPL favors mid-range examples to avoid trivial or outlier data. These scores are cheap to compute and have shown to improve sample efficiency in the instruction tuning setting (yin2025computeconstraineddataselection; ankner2024perplexedperplexityperplexitybaseddata).
@@
Embedding-based methods use data representations to assign scores. S1 (s1) enforces coverage by classifying problems into topics with an LLM and sampling across them. RDS+ (rds) computes pairwise scores $s(v,d)=\langle E(v),E(d)\rangle$ between validation and training examples, selecting examples most relavent to each $v\in\mathcal{V}$ in a round-robin fashion.
@@
Gradient and influence based methods compute $s(v,d)$ by estimating the causal contribution of training examples on validation loss. TracIn (tracin) sums gradient inner products across training checkpoints and LESS (LESS) scales this approach to large models using gradient datastores. InfDist (infdist) learns optimal per-example weights so that a single gradient step minimizes validation loss.
@@
Unlike gradient-similarity methods that track training dynamics, influence functions measure the effect of a training example on converged model parameters and downstream objectives $f(\theta^{\star})$. Formally, the influence of a training example $d\in\mathcal{D}$ is $\mathcal{I}_{f}(d)=-\nabla_{\theta}f(\theta^{\star})^{\top}H^{-1}\nabla_{\theta}L(d,\theta^{\star})$ where $H$ is the Hessian of the training objective at $\theta^{\star}$. Approximations such as EK-FAC have made this tractable for LLMs (grosse2023studyinglargelanguagemodel). ruis2025proceduralknowledgepretrainingdrives shows that examples with high estimated influence on cross-entropy loss also causally affect downstream accuracy, and their removal causes larger performance drops than random pruning or TracIn.
@@ 3.1 Influence Functions Scoring
Let $\theta$ be a pretrained model and $\mathcal{D}$ the LIMO training dataset (ye2025limoreasoning). We fine-tune $\theta$ on $\mathcal{D}$ for 10 epochs following the settings of ye2025limoreasoning, obtaining $\theta^{\text{LIMO}}$. We then evaluate on the MATH500 benchmark $\mathcal{V}$ for both model parameters $\theta$ and $\theta^{\text{LIMO}}$ (hendrycks2021measuring).
@@
Our goal is to estimate the causal influence of training examples $d\in\mathcal{D}$ on these outcome sets. Because influence functions require a differentiable proxy, we follow grosse2023studyinglargelanguagemodel and compute influence with respect to cross-entropy loss on completions. For each validation query $v\in C\cup I$ with completion $y_{c}$, the influence of a training point $d$ is:
@@ 4 Experiments and Results
We perform supervised fine-tuning (SFT) on pruned versions of the LIMO dataset, using two models: LLaMA-3-8B-Instruct (grattafiori2024llama), also used during pruning, and Qwen2.5-Math-7B-Instruct (yang2024qwen2), external to the pruning process. We ran full finetuning following the training settings from LIMO for 10 epochs, saving checkpoints at every other epoch. Given that we already were optimizing for MATH500, we treated it as a validation set and chose the checkpoint which performed best on MATH500. We chose this strategy as for different datasets, optimal checkpoints were not consistent.
Evaluation focuses on the math reasoning benchmarks: AIME24, AMC23, OlympiadBench (he2024olymbpiadbench), GSM8k (cobbe2021training). Additionally we include GPQA (rein2024gpqa) to assess transfer to non-math reasoning. We compare three IF-based pruning strategies (Correct, Incorrect, Combined) against Random, Mid-PPL , and RDS+ (rds), and report also the unpruned dataset and the base model for reference. All results use the unbiased pass@1 metric (chen2021evaluating) with N=8N{=}8 samples in a zero-shot chain-of-thought setting, following the LIMO protocol. Full training and evaluation details are provided in the Appendix.
@@ Appendix A Training and Evaluation Details
We follow the evaluation setup of LIMO (ye2025limoreasoning), with the addition of the GSM8k test set manually added. We report the unbiased pass@1 metric (chen2021evaluating), estimated from N=8N{=}8 sampled generations. Sampling uses temperature $0.6$, top-$p=0.95$, top-$k=1$, and a maximum generation length of 16,38416{,}384 tokens.
@@ Appendix B Notes on influence functions computation
To compute influence function scores, we follow the derivation and approximations described in grosse2023studyinglargelanguagemodel. Following ruis2025proceduralknowledgepretrainingdrives,we perform the computation considering only MLP layers.